Find the integral surface of the partial differential equation
(X-y)y^2p+(y-x)x^2q=(x^2+y^2)z
1
Expert's answer
2020-03-18T17:16:03-0400

dx/((x-y)y2)=dy/((y-x)x2)=dz/((x2+y2)z)

dx/((x-y)y2)=dy/((y-x)x2)

"\int" y2dy = "\int" -x2dx

y3 = -x3+c1

dx/((x-y)y2)=dz/((x2+y2)z)

"\int" y2/z dz = "\int" (x2+y2)/(x-y) dx


"\intop" (x2+y2)/(x-y) dx = "\int" ((2y2/(x-y)+x+y)dx = 2y2"\intop" 1/(x-y)dx +"\intop" x dx + y "\intop" 1 dx =

= 2 y2 ln(x-y) + x2/2 + xy + c2


y2ln(z) = 2 y2 ln(x-y) + x2/2 + xy + c2

integral surface:

y3 = -x3+c1

y2ln(z) = 2 y2 ln(x-y) + x2/2 + xy + c2